OpenStudy (itsme.caro):

lim ∂->∞ ∂^2/(1-cos∂)

OpenStudy (irishboy123):

you killed the previous question too quickly. you can't reopen it but you established that \(\lim\limits_{x \to \infty} \cos \left( \frac{1}{x} \right)\) \(= \cos \left( \lim\limits_{x \to \infty} \frac{1}{x} \right) = \cos 0 = 1\), because cosine is a continuous function you can use that here

OpenStudy (idku):

\(\color{black}{\displaystyle \lim _{x\to \infty }\frac{x^2}{1-\cos x}=\infty. }\) You know that \(\color{black}{\displaystyle \cos x }\) is at least \(\color{black}{\displaystyle -1 }\) and at most \(\color{black}{\displaystyle 1 }\). When \(\color{black}{\displaystyle \cos x=1 }\), you have (so to speak) \(\color{black}{\displaystyle \infty^2/0=\infty^3=\infty }\) (Well, not like that, but you know it then diverges to \(\color{black}{\displaystyle \infty}\).) When \(\color{black}{\displaystyle \cos x }\) is any value \(\color{black}{\displaystyle \varepsilon }\), such that \(\color{black}{\displaystyle \{\varepsilon:\varepsilon\in[-1,1]~~{\rm and}~~\varepsilon\ne0\}}\) (I've already excluded \(\color{black}{\displaystyle \varepsilon=0 }\) by excluding \(\color{black}{\displaystyle \cos x=1 }\)) \(\color{black}{\displaystyle \bf THEN{\tiny~}}\) you have the following: \(\color{black}{\displaystyle \lim _{x\to \infty }\frac{x^2}{\varepsilon}=\frac{1}{\varepsilon} \lim _{x\to \infty }x^2=\infty}\).

OpenStudy (idku):

This is all just common sense, really ... all I did is just speak it out.

OpenStudy (idku):

A more interesting case though would be when \(\color{black}{\displaystyle x\to0}\). \(\color{black}{\displaystyle \lim _{x\to 0 }\frac{x^2}{1-\cos x}=\lim _{x\to 0 }\frac{2x}{\sin x}=\lim _{x\to 0 }\frac{2}{\cos x}=\frac{2}{\cos(0)}=1}\) By, LHS twice.